12. A graduation cap is thrown straight up from the ground with an unknown velocity. It reaches its highest point after 12 s. With what Initial velocity did it leave the ground? Formula: v0 = v -gt


13.Using his elastic powers, Luffy launches himself forward, from rest, with an acceleration of 11 m/ If he covered 24 m, what was his final velocity?

Formula: v = 2ax


14. Doc Ock scales a bridge with an acceleration of 2.5 m/s'. As a result, the accelerates from rest to 8.6 m/s. How much distance did he cover?

Formula: v = 2ax

15. A Honda Civic accelerates from rest to 26 m/s over a distance of 65 m. What acceleration experience?

Formula: v= 2ax



1
Expert's answer
2022-01-16T13:08:53-0500

"12.\space v_0=gt=9.8·12=117.6\text{ m/s}.\\
13.\space v=\sqrt{2ax}=22.98\text{ m/s}.\\
14.\space x=v^2/(2a)=14.79\text{ m}.\\
15.\space a=v^2/(2x)=5.2\text{ m/s}^2."